The dimensions of a brick in inches are 2 by 4 by 9. How many such bricks are needed to have a total volume of exactly 2 cubic feet?
Answer by checkley77(12844)

2*4*9=72 IN^3
A=S^3
2=S^3
S=CUBERT2
S=1.26 FT.
1.26*12=15.12 INCHES.
15.12^3=3456.65 IN^3
3456.65/72=48 BRICKS.